OpenStudy (anonymous):

Write the slope-intercept inequality for the graph below. If necessary, use <= for or >= for

Vocaloid (vocaloid):

picture of the graph, please?

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

@Redneckk__Lifee @Vocaloid

Vocaloid (vocaloid):

ok, let's "ignore" the shading for now and just focus on the line do you know how to write an equation for the line, in point-intercept form?

OpenStudy (anonymous):

to be honest no

Vocaloid (vocaloid):

no problem, first step is to use the slope equation (x1,y1) = (0,1) (x2,y2) = (3,-1) slope = m = (y2-y1)/(x2-x1) = ?